# Modulo arithmetic of function spaces: Subset hyperspaces as quotients of function spaces

### v1

@@ Abstract.
Let $X$ be a (topological) space and $Cl(X)$ the collection of nonempty closed subsets of $X$. Given a topology on $Cl(X)$, making $Cl(X)$ a space, a (subset) hyperspace of $X$ is a subspace $\mathcal{J}\subset Cl(X)$ with an embedding $X\hookrightarrow\mathcal{J}$, $x\mapsto\{x\}$ (which thus requires $X$ to be $T_{1}$). In this note, we characterize certain hyperspaces $\mathcal{J}\subset Cl(X)$ as explicit quotient spaces of function spaces $\mathcal{F}\subset X^{Y}$ and discuss metrization of associated compact-subset hyperspaces in this setting.
@@
Hyperspace, function space, quotient topology, swrc-topology.
@@ 1. Introduction
(subset space) = (function space) modulo (quotient map).
Remark 3.18 (5)-(7) makes a connection with quotients of groups, rings, and modules.
@@
Given a set $X$ and a collection $\mathcal{B}\subset\mathcal{P}(X)$ of subsets of $X$, the topology on $X$ “generated by $\mathcal{B}$”, i.e., “with $\mathcal{B}$ as a subbase”, is denoted by $\langle\mathcal{B}\rangle$. Let $X=(X,\tau)$ be a (topological) space. A topology $\tau^{\prime}\subset\mathcal{P}(X)$ on $X$ is $\tau$-compatible if $\tau\subset\tau^{\prime}$ (where $\tau^{\prime}$ can therefore inherit certain desirable properties of $\tau$, e.g., if $\tau$ is Hausdorff, regular, or normal respectively, then so is $\tau^{\prime}$). If $A\subset X$, the closure of $A$ in $X$ is denoted by $cl_{X}(A)$, or by $\overline{A}$ if the underlying space $X$ is understood. We say that $A\subset X$ is precompact in $X$ if $cl_{X}(A)$ is compact in $X$. Let $Z$ be a metric space. The completion of $Z$ is (up to isometry) the complete metric space $\widetilde{Z}$ containing $Z$ as a dense subset. We say $A\subset X$ is totally bounded if $A$ is precompact in $\widetilde{Z}$ (and hence also simply called “precompact” when the context is understood).
@@
Let $\mathcal{H}(X)$ denote the class of homeomorphisms of $X$ and let $\mathcal{C}\subset\mathcal{H}(X)$. By geometry of $X$ (resp., $\mathcal{C}$-geometry of $X$) we mean the study of one or more properties of $X$ that are invariant under specific homeomorphisms of $X$ (resp., homeomorphisms of $X$ in $\mathcal{C}$), where the invariant properties are accordingly called geometric properties (resp., $\mathcal{C}$-geometric properties) of $X$. By metric geometry we mean geometry that employs metrics. With $\mathcal{P}^{\ast}(X)$ denoting the set of nonempty subsets of $X$, we would like to study the simplest kinds of topologies on $\mathcal{J}\subset\mathcal{P}^{\ast}(X)$ using better understood topologies on $\mathcal{F}\subset X^{Y}$ (for sets $Y$). Since $X$ can be seen as a subset of $\mathcal{P}^{\ast}(X)$ in a natural way (through the inclusion $X\hookrightarrow\mathcal{P}^{\ast}(X),~{}x\mapsto\{x\}$), we often consider those topologies on $\mathcal{J}\supset\textrm{Singlt}(X)$,
$\textrm{Singlt}(X):=\big{\{}\{x\}:x\in X\big{\}},$
that can be seen as extensions of the topology of $X$. Such topologies are called hypertopologies of $X$, and the associated spaces $\mathcal{J}$, $\textrm{Singl}(X)\subset\mathcal{J}\subset\mathcal{P}^{\ast}(X)$, are called (subset) hyperspaces of $X$. A closed-subset hyperspace (resp. compact-subset hyperspace, resp. bounded-subset hyperspace) of $X$ is a hyperspace consisting of closed subsets (resp. compact subsets, resp. bounded subsets) of $X$.
@@
Let $X$ be a $T_{1}$ space, $Cl(X)\subset\mathcal{P}^{\ast}(X)$ the set of nonempty closed subsets of $X$ as a hyperspace whose topology is comparable to the Vietoris topology $\tau_{v}$, $K(X)\subset Cl(X)$ the subspace consisting of all nonempty compact subsets of $X$, and $FS_{n}(X):=\{A\in K(X):1\leq|A|\leq n\}$ the subspace consisting of all nonempty $n$-finite subsets of $X$. When $X$ is a metric space, we further let $BCl(X)\subset Cl(X)$ denote the subspace consisting of all nonempty bounded closed subsets of $X$. Let $Y$ be a set. On relevant sets of functions $\mathcal{F}\subset X^{Y}$, besides the standard product topology (or topology of pointwise convergence) $\tau_{p}$ and related topologies, we will introduce a preferred topology $\tau_{\pi}$ on $\mathcal{F}\subset X^{Y}$ (Definition 3.7) in order to discuss metrization of compact-subset hyperspaces (Theorems 4.8,4.12). Since a closed set $C\in Cl(X)$ can be seen as the closure $cl_{X}(f(Y))$, or $\overline{f(Y)}$, of the image $f(Y)$ of some function $f\in X^{Y}$, we will describe specific hyperspaces $\mathcal{J}\subset Cl(X)$ as explicit $\tau_{\pi}$-quotient spaces of function spaces $(\mathcal{F},\tau_{\pi})$ using the image-closure assignment (named unordering map in Definition 3.1)
$q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau_{\pi q}),~{}f\mapsto cl_{X}(f(Y))$
@@
as the quotient map, where $\tau_{\pi q}$ is the $\tau_{\pi}$-quotient topology (footnote111 That is, $\tau_{\pi q}:=\sup\big{\{}\tau~{}|~{}q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau)~{}\textrm{is surjective and continuous}\big{\}}$ is the largest topology on $Cl_{Y}(X)$ with respect to which the map $q:(\mathcal{F},\tau_{\pi})\rightarrow Cl_{Y}(X)$ is surjective and continuous. ) on $Cl_{Y}(X):=q(\mathcal{F})$.
@@
First, we would like to have a straightforward description (say by means of a natural quotient map) of the relation between basic function space topologies (e.g., the topologies of pointwise convergence, compact convergence, uniform convergence) and the Vietoris topology of subset hyperspaces. Second, if $X$ is a metric space, then as indicated in Question 6.3, we seek a suitable geometric framework for answering [akofor2024, Question 5.1] concerning characterization/representation of paths in $BCl(X)$ in terms of paths in $X$ or $\widetilde{X}$. This appears to require a deeper understanding of the relationship between basic function space topologies and the metrization of associated hypertopologies by the Hausdorff distance $d_{H}$ (Equation (2.1)).
Our work is therefore primarily motivated by the need to use metrizability of certain hyperspaces $\mathcal{J}\subset K(X)$, viewed as quotients of function spaces $\mathcal{F}\subset X^{Y}$ (suitably topologized), say as in [bartsch2014], to gain a deeper understanding/interpretation of $d_{H}$ as a metric on $\mathcal{J}$. Our main results (Theorems 4.4,4.8,4.12,5.4) are based on the observation (through Lemmas 3.14,3.16 and Theorem 3.17) that basic/$\tau_{v}$-compatible hyperspace topologies indeed seem natural to describe as quotients of basic/$\tau_{p}$-compatible function space topologies.
Summary and highlights of main results
To describe our main results, we consider the following question.
  1. Question A: Let $X$ be a space, $Y$ a set, $\mathcal{F}\subset X^{Y}$ a subset such that $q(\mathcal{F})=q(X^{Y})$, $Cl_{Y}(X):=q(X^{Y})$, and $K_{Y}(X):=Cl_{Y}(X)\cap K(X)$, where $q$ is the map given by
@@
    If $\tau\supset\tau_{v}$ is a topology on $Cl_{Y}(X)$, does there exist a topology $\widetilde{\tau}\supset\tau_{p}$ on $\mathcal{F}$ such that $\tau_{v}\subset\widetilde{\tau}_{q}\subset\tau$ (where $\widetilde{\tau}_{q}$ is the quotient topology induced on $Cl_{Y}(X)$ by $\widetilde{q}$ through the map $q$)?
In Section 3, we describe a class of hyperspaces as quotients of function spaces, noting that some related work has been considered in [bartsch2014] (see Fact 1 below). Next in Section 4, we discuss the concrete realization of certain preferred function space topologies and metrization of compact-subset hyperspaces, and (under some regularity constraints on $\mathcal{F}$) give a positive answer to Question A (1a)&\&&(1b) in Theorems 4.4,4.12 respectively. In Section 5 we discuss the realization of $\tau_{v}$-compatible hyperspace topologies as quotients of $\tau_{p}$-compatible function space topologies, and give a positive answer to Question A(2) in Theorem 5.4.
Fact 1 (Remark 3.10 (1)).
@@
$q:C(Y,X)\subset(X^{Y},\tau_{\textrm{co}})\rightarrow(K(X),\tau_{v}),~{}~{}f\mapsto cl_{X}(f(Y)),$
where $C(Y,X):=\{\textrm{continuous}~{}f\in X^{Y}\}$ and $\tau_{co}$ is the compact-open topology (Definition 3.9).
@@
Throughout, we appeal to intuition by preferably employing sequences and nets (instead of open sets) in our concepts, results, and proofs whenever this seems convenient.
@@ 2. Preliminary remarks
One of our main results of this section, namely, metrization of compact-subset hyperspaces in Theorem 4.12, is well-known for standard hyperspaces with the Vietoris topology (see Lemma 4.7, [IllanNadl1999, Theorem 2.4, Theorem 3.1, Theorems 3.2,3.4], [michael1951, Theorem 4.9.13], and [DolMyn2010, Bart2014, McCoy1998, RodRom2002, LeviEtal1993, Render1993]). Our goal here is simply to present a specialized review that emphasizes the description of $\tau_{v}$-compatible hyperspaces as quotients of $\tau_{p}$-compatible function spaces (where the role originally played by the Vietoris topology $\tau_{v}$ is now played by $\tau_{v}$-compatible quotient topologies induced by relevant function space topologies). The discussion may be viewed as an extension of preliminary discussions for the case of finite-subset hyperspaces in [akofor2020, Chapter 1, especially around Definition 1.0.1 and Proposition 1.2.2].
Definition 2.1 (Hausdorff distance).
Let $X$ be a metric space $X$, $x\in X$, $A\subset X$, $\varepsilon>0$,
$\textstyle d(x,A)=\operatorname{dist}(x,A):=\inf\limits_{a\in A}d(x,a),$
$\textstyle N_{\varepsilon}(A):=\{x\in X:\operatorname{dist}(x,A)<\varepsilon\},~{}\textrm{and}~{}$
$\overline{N}_{\varepsilon}(A)=A_{\varepsilon}:=\{x\in X:\operatorname{dist}(x,A)\leq\varepsilon\}.$
@@
(2.1) $d_{H}(A,B)$ $\textstyle:=\max\{\sup_{a\in A}\operatorname{dist}(a,B),\sup_{b\in B}\operatorname{dist}(b,A)\}$
@@
$=\textstyle\sup_{x\in A\cup B}|\operatorname{dist}(x,A)-\operatorname{dist}(x,B)|$
$\textstyle=\sup_{x\in X}|\operatorname{dist}(x,A)-\operatorname{dist}(x,B)|.$
Definition 2.2 (Quotient map, Quotient space, Quotient topology, Saturated set).
Let $U=(U,\tau_{U})$ and $V=(V,\tau_{V})$ be spaces and $f:U\rightarrow V$ a map. Then $f$ is called a quotient map if (i) $f$ is surjective and (ii) a set $B\subset V$ is open if and only if its preimage $f^{-1}(B)\subset U$ is open. If $f$ is a quotient map, we say $V$ is a quotient space (or a quotient) of $U$ and $\tau_{V}$ is the $f$-quotient topology induced on $V$ by $\tau_{U}$, written as
[…]
@@
Remark 2.3.
@@
    If $I\subset U$, the restriction $q|_{I}:I\rightarrow q(I)$ need not be a quotient map with respect to the subspace topologies.
@@
    If $O\subset U$ (resp., $C\subset U$) is a $q$-saturated open (resp., closed) set, then $q|_{O}:O\rightarrow q(O)$ (resp., $q|_{C}:C\rightarrow q(C)$) is a quotient map with respect to the subspace topologies. Indeed, if $I\subset U$ satisfies $q^{-1}q(I)=I$, then for any $B\subset V$,
    $q|_{I}^{-1}(q(I)\cap B)$ $:=I\cap q^{-1}(q(I)\cap B)=q^{-1}q(I)\cap q^{-1}(q(I)\cap B)=q^{-1}\big{(}q(I)\cap q(I)\cap B\big{)}$
    $=q^{-1}\big{(}q(I)\cap B\big{)}.$
@@
    In particular, if $O\subset V$ (resp., $C\subset V$) is an open (resp., closed) set, then $q^{-1}(O)\subset U$ (resp., $q^{-1}(C)\subset U$) is a $q$-saturated open (resp. closed) set, giving respective quotient maps
    $q|_{q^{-1}(O)}:q^{-1}(O)\rightarrow O,~{}~{}~{}~{}q|_{q^{-1}(C)}:q^{-1}(C)\rightarrow C.$
@@
which need not be restrictions of quotient maps $(X^{Y},\tau^{\prime})\rightarrow(q(X^{Y}),\tau^{\prime}_{q})$, even if $(\mathcal{F},\tau)\subset(X^{Y},\tau)$ (i.e., the topology of $\mathcal{F}$ happens to be a subspace topology of a topology on $X^{Y}$). (footnote333 If $A$ is a set and $(B,\tau)$ a space such that $A\subset B$, then the subspace topology on $A$ is the topology $\tau^{i}$ that makes the inclusion $i:(A,\tau^{i})\hookrightarrow(B,\tau)$ a quotient map, where we know $\tau^{i}=\tau\cap A:=\{U\cap A:U\in\tau\}$. Similarly, if $(A,\tau)$ is a space and $B$ a set such that $A\subset B$, then the inclusion $i:A\hookrightarrow B$ induces the $i$-quotient topology (call it the superspace topology) $\tau_{i}$ on $B$ given by $\tau_{i}=\{U\subset B:U\cap A\in\tau\}$. )

### v2

@@ Abstract.
Let $X$ be a (topological) space and $Cl(X)$ the collection of nonempty closed subsets of $X$. Given a topology on $Cl(X)$, making $Cl(X)$ a space, a (subset) hyperspace of $X$ is a subspace $\mathcal{J}\subset Cl(X)$ with an embedding $X\hookrightarrow\mathcal{J}$, $x\mapsto\{x\}$. In this note, we characterize certain hyperspaces $\mathcal{J}\subset Cl(X)$ as explicit quotient spaces of function spaces $\mathcal{F}\subset X^{Y}$ and discuss metrization of associated compact-subset hyperspaces in this setting. In particular, we find that any hyperspace topology containing the Vietoris topology is a quotient of a function space topology containing the topology of pointwise convergence.
@@
Subset hyperspace, function space, quotient topology, swrc-topology.
@@ 1. Introduction
(subset hyperspace) = (function space) modulo (quotient map).
Theorem 3.3’s Remark (5)-(7) makes a connection with quotients of groups, rings, and modules.
@@
Given a set $X$ and a collection $\mathcal{B}\subset\mathcal{P}(X)$ of subsets of $X$, the topology on $X$ generated by $\mathcal{B}$, i.e., with $\mathcal{B}$ as a subbase, is denoted by $\langle\mathcal{B}\rangle$. Given two (topological) spaces $X_{1}$ and $X_{2}$, $X_{1}\cong X_{2}$ means $X_{1}$ and $X_{2}$ are homeomorphic. Let $X=(X,\tau)$ be a space. A topology $\tau^{\prime}\subset\mathcal{P}(X)$ on $X$ is $\tau$-compatible if $\tau\subset\tau^{\prime}$ (where $\tau^{\prime}$ can therefore inherit certain desirable properties of $\tau$, e.g., if $\tau$ is $T_{0}$, $T_{1}$, or $T_{2}$ respectively, then so is $\tau^{\prime}$). If $A\subset X$, the closure of $A$ in $X$ is denoted by $cl_{X}(A)$, or by $\overline{A}$ if the underlying space $X$ is understood. We say that $A\subset X$ is precompact in $X$ if $cl_{X}(A)$ is compact in $X$. Let $Z$ be a metric space. The completion of $Z$ is (up to isometry) the complete metric space $\widetilde{Z}$ containing $Z$ as a dense subset. We say $A\subset Z$ is totally bounded if $A$ is precompact in $\widetilde{Z}$. It is easy to verify that a set $A\subset Z$ is totally bounded in the usual sense if and only if $A$ is precompact in $\widetilde{Z}$. It might be worth noting that in the context of [akofor2024], “precompact” is automatically equivalent to “totally bounded”, as explained by [akofor2024, Notation 1].
@@
Let $\mathcal{H}(X)$ denote the class of homeomorphisms of $X$ and let $\mathcal{C}\subset\mathcal{H}(X)$. By geometry of $X$ (resp., $\mathcal{C}$-geometry of $X$) we mean the study of one or more properties of $X$ that are invariant under specific homeomorphisms of $X$ (resp., homeomorphisms of $X$ in $\mathcal{C}$), where the invariant properties are accordingly called geometric properties (resp., $\mathcal{C}$-geometric properties) of $X$. By metric geometry we mean geometry that employs metrics. With $\mathcal{P}^{\ast}(X)$ denoting the set of nonempty subsets of $X$, we would like to study the simplest kinds of topologies on $\mathcal{J}\subset\mathcal{P}^{\ast}(X)$ using better understood topologies on $\mathcal{F}\subset X^{Y}$ (for sets $Y$). Since a $T_{1}$ space $X$ can be seen as a subset of $\mathcal{P}^{\ast}(X)$ in a natural way (through the inclusion $X\hookrightarrow\mathcal{P}^{\ast}(X),~x\mapsto\{x\}$), we often consider those topologies on $\mathcal{J}\supset\textrm{Singlt}(X):=\big\{\{x\}:x\in X\big\}$ that can be seen as extensions of the topology of $X$. Such topologies are called hypertopologies of $X$, and the associated spaces $\mathcal{J}$, $\textrm{Singlt}(X)\subset\mathcal{J}\subset\mathcal{P}^{\ast}(X)$, are called (subset) hyperspaces of $X$. A closed-subset hyperspace, compact-subset hyperspace, or bounded-subset hyperspace of $X$ is a hyperspace consisting respectively of closed subsets, compact subsets, or bounded subsets of $X$.
@@
Let $X$ be a $T_{1}$ space, $Cl(X)\subset\mathcal{P}^{\ast}(X)$ the set of nonempty closed subsets of $X$ as a hyperspace whose topology is comparable to the Vietoris topology $\tau_{v}$, $K(X)\subset Cl(X)$ the subspace consisting of all nonempty compact subsets of $X$, and $FS_{n}(X):=\{A\in K(X):1\leq|A|\leq n\}$ the subspace consisting of all nonempty finite subsets of $X$ of cardinality at most $n$. When $X$ is a metric space, we further let $BCl(X)\subset Cl(X)$ denote the subspace consisting of all nonempty bounded closed subsets of $X$. Let $Y$ be a set. On relevant sets of functions $\mathcal{F}\subset X^{Y}$, besides the standard product topology (or topology of pointwise convergence) $\tau_{p}$ and related topologies, we will introduce a preferred topology $\tau_{\pi}$ on $\mathcal{F}\subset X^{Y}$ (Definition 7) in order to discuss metrization of compact-subset hyperspaces (Theorems 4.6 and 4.9). Since a closed set $C\in Cl(X)$ can be seen as the closure of the image $f(Y)$ of some function $f\in X^{Y}$, the image-closure assignment (named unordering map in Definition 3) given by
(1) $q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau_{\pi q}),~f\mapsto cl_{X}(f(Y))$
@@
induces a quotient topology $\tau_{\pi q}$, the $\tau_{\pi}$-quotient topology (footnote [^1] That is, $\tau_{\pi q}:=\sup\big\{\tau~|~q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau)~\textrm{is surjective and continuous}\big\}$ is the largest topology on $Cl_{Y}(X)$ with respect to which the map $q:(\mathcal{F},\tau_{\pi})\rightarrow Cl_{Y}(X)$ is surjective and continuous. ), on the $Y$-indexed closed subsets
(2) $Cl_{Y}(X):=q(\mathcal{F}).$
@@
First, we would like to have a straightforward description (by means of a natural quotient map) of the relation between basic function space topologies (e.g., the topologies of pointwise convergence, compact convergence, uniform convergence) and the Vietoris topology of subset hyperspaces. We expect this to facilitate the use of function spaces to study subset hyperspaces and vice versa.
Second, for a metric space $X$, we seek a suitable geometric framework for answering [akofor2024, Question 5.1] concerning characterization/representation of Lipschitz paths in $BCl(X)$ in terms of Lipschitz paths in $X$ or $\widetilde{X}$ (as indicated in Question 5). Due to the key role of Hausdorff distance in the results of [akofor2024], the desired geometric framework appears to require a deeper understanding of the relationship between certain basic function space topologies and the metrization of associated hypertopologies by the Hausdorff distance $d_{H}$ (Equation (3)).
Our work is therefore primarily motivated by the need to use metrizability of certain hyperspaces $\mathcal{J}\subset K(X)$, viewed as quotients of function spaces $\mathcal{F}\subset X^{Y}$ (suitably topologized), to strengthen our understanding/interpretation of $d_{H}$ as a key metric on $\mathcal{J}$. Our main results (Theorems 4.2, 4.6, 4.9, and 5.1) are based on the observation (via Lemma 3.1, Lemma 3.2, and Theorem 3.3) that basic hyperspace topologies (with reasonable separation properties) admit a natural description as quotients of basic function space topologies (with reasonable separation properties).
Summary and highlight of main results
To describe our main results, consider the following question.
  1. Question A: Let $X$ be a space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$ a subset satisfying $q(\mathcal{F})=q(X^{Y})$. Define $Cl_{Y}(X):=q(X^{Y})$ and $K_{Y}(X):=Cl_{Y}(X)\cap K(X)$, where $q$ is the map given by
@@
    If $\tau\supset\tau_{v}$ is a topology on $Cl_{Y}(X)$, does there exist a topology $\widetilde{\tau}\supset\tau_{p}$ on $\mathcal{F}$ such that $\tau\supset\widetilde{\tau}_{q}\supset\tau_{v}$ (where $\widetilde{\tau}_{q}$ is the quotient topology induced on $Cl_{Y}(X)$ by $\widetilde{\tau}$ via the map $q$)?
In Section 3, we describe a class of hyperspaces as quotients of function spaces, noting that some related work has been considered in [bartsch2014] (see Fact 1 below). Next in Section 4, we discuss the concrete realization of certain preferred function space topologies and metrization of compact-subset hyperspaces, and (under some regularity constraints on $\mathcal{F}$) give a positive answer to Question A (1a)&\&(1b) in Theorems 4.2 &\& 4.9 respectively. In Section 5, we discuss the realization of $\tau_{v}$-compatible hyperspace topologies as quotients of $\tau_{p}$-compatible function space topologies, and give a positive answer to Question A(2) in Theorem 5.1.
Fact 1 (Definition 8’s Remark (1)).
@@
$q:C(Y,X)\subset(X^{Y},\tau_{\textrm{co}})\rightarrow(K(X),\tau_{v}),$
where $C(Y,X):=\{\textrm{continuous}~f\in X^{Y}\}$ and $\tau_{co}$ is the compact-open topology (Definition 8).
@@
Throughout, we appeal to intuition by preferably employing sequences and nets (instead of open sets) in our results and proofs whenever this seems convenient.
@@ 2. Preliminary remarks
One of our main results, namely, metrization of compact-subset hyperspaces in Theorem 4.9, is well-known for standard hyperspaces with the Vietoris topology (see Lemma 4.5, [IllanNadl1999, Theorems 2.4, 3.1, 3.2, 3.4], [michael1951, Theorem 4.9.13], and [DolMyn2010, Bart2014, McCoy1998, RodRom2002, LeviEtal1993, Render1993]). Our goal here is simply to present a specialized review that emphasizes the description of $\tau_{v}$-compatible subset hyperspaces as quotients of $\tau_{p}$-compatible function spaces (where the role originally played by the Vietoris topology $\tau_{v}$ is now played by $\tau_{v}$-compatible quotient topologies induced by relevant function space topologies). The discussion may be viewed as an extension of preliminary discussions for the case of finite-subset hyperspaces in [akofor2020, Chapter 1, especially around Definition 1.0.1 and Proposition 1.2.2].
Definition 1 (Hausdorff distance).
Let $X$ be a metric space, $x\in X$, $A\subset X$, $\varepsilon>0$,
$\textstyle d(x,A):=\inf\limits_{a\in A}d(x,a),$
$\textstyle N_{\varepsilon}(A):=\{x\in X:d(x,A)<\varepsilon\},~\textrm{and}$~
$\overline{N}_{\varepsilon}(A)=A_{\varepsilon}:=\{x\in X:d(x,A)\leq\varepsilon\}.$
@@
(3) $d_{H}(A,B)$ $\textstyle:=\max\{\sup_{a\in A}d(a,B),\sup_{b\in B}d(b,A)\}$
@@
$=\textstyle\sup_{x\in A\cup B}|d(x,A)-d(x,B)|$
$\textstyle=\sup_{x\in X}|d(x,A)-d(x,B)|.$
Definition 2 (Saturated set).
@@
Recall that a quotient map $f:(U,\tau_{U})\rightarrow(V,\tau_{V})$ makes $V$ a quotient (or quotient space) of $U$, in which case we will call $\tau_{V}$ the $f$-quotient topology induced on $V$ by $\tau_{U}$, written as
$\tau_{V}=\tau_{Uf},$
which is a prelude to the notation introduced in Item 4 of Definition 3.
Remark.
@@
    Given $I\subset U$, the restriction $q|_{I}:I\rightarrow q(I)$ need not be a quotient map with respect to the subspace topologies. Indeed, if $P$ is open in $V$, i.e., $E:=P\cap q(I)$ is open in $q(I)$, then $q|_{I}^{-1}(E):=q^{-1}(E)\cap I=q^{-1}(P)\cap I$ (which is open in $I$), but $q^{-1}(F)\cap I=q^{-1}(P)\cap I$ (for some $F\subset V$) does not always imply that $F\cap q(I)$ is open in $q(I)$.
@@
    If $O\subset U$ (resp., $C\subset U$) is a $q$-saturated open (resp., closed) set, then $q|_{O}:O\rightarrow q(O)$ (resp., $q|_{C}:C\rightarrow q(C)$) is a quotient map with respect to the subspace topologies. Indeed, in Part (1) above with $I$ replaced by $O$, $q^{-1}(F)\cap O=q^{-1}(P)\cap O$ ($\iff$ $F\cap q(O)=P\cap q(O)$) is open in $U$ if and only if $F\cap q(O)$ is open in $q(O)$.
@@
    In particular, if $O\subset V$ (resp., $C\subset V$) is an open (resp., closed) set, then $q^{-1}(O)\subset U$ (resp., $q^{-1}(C)\subset U$) is a $q$-saturated open (resp. closed) set, giving a quotient map
    $q|_{q^{-1}(O)}:q^{-1}(O)\rightarrow O,~~(\textrm{resp.,}~~q|_{q^{-1}(C)}:q^{-1}(C)\rightarrow C).$
@@
which need not be restrictions of quotient maps $(X^{Y},\tau^{\prime})\rightarrow(q(X^{Y}),\tau^{\prime}_{q})$, even if $(\mathcal{F},\tau)\subset(X^{Y},\tau)$ (i.e., the topology of $\mathcal{F}$ happens to be a subspace topology of a topology on $X^{Y}$). (footnote [^2] If AA is a set and $(B,\tau)$ a space such that $A\subset B$, then the subspace topology on $A$ is the topology $\tau^{i}$ that makes the inclusion $i:(A,\tau^{i})\hookrightarrow(B,\tau)$ a quotient map, where we know $\tau^{i}=\tau\cap A:=\{U\cap A:U\in\tau\}$. Similarly, if $(A,\tau)$ is a space and $B$ a set such that $A\subset B$, then the inclusion $i:A\hookrightarrow B$ induces the $i$-quotient topology (call it the superspace topology) $\tau_{i}$ on $B$ given by $\tau_{i}=\{U\subset B:U\cap A\in\tau\}$. )